v(t)= A (1-e^(-t/tmaxspeed))
1.Identify the
●units of the coefficient A
●physical meaning of A
●velocity of the car at t = 0
●Asymptote of this function as t → ∞?
2.Sketch a graph of velocity vs. time.
3.Derive an equation x(t) for the instantaneous position of the car as a function of time. Identify the
●value x when t = 0 s
●asymptote of this function as t → ∞
4.Sketch a graph of position vs. time.
5.Derive an equation for the instantaneous acceleration of the car as a function of time. Identify the
●acceleration of the car at t = 0 s
●asymptote of this function as t → ∞
6.Sketch a graph of acceleration vs. time.
7.Apply your mathematical models to your allocated car. Use the given data for the 0 – 28 m/s and 400m times to calculate the:
●value of the coefficient A
●maximum velocity
Maximum acceleration.
1
Expert's answer
2020-05-01T18:48:19-0400

QUESTION 1


Since we know exactly the dimensions of some expressions that are included in the formula, we can conclude



"\left\{\begin{array}{l}
[v(t)]=\left[\displaystyle\frac{m}{sec}\right]\\[0.3cm]
\left[1-e^{-t/t_{maxspeed}}\right]=[\text{just a number}]
\end{array}\right.\rightarrow [A]=\left[\frac{m}{sec}\right]"

We substitute "t=0" and find the value of velocity :



"v(0)=A\cdot\left(1-e^{-\displaystyle\frac{0}{t_{maxspeed}}}\right)=A\cdot(1-1)=0\\[0.3cm]
\boxed{v(0)=0}"

To find the asymptote as "t\to+\infty" we calculate the limit :



"v_{\infty}=\lim\limits_{t\to+\infty}\left(A\cdot\left(1-e^{-\displaystyle\frac{t}{t_{maxspeed}}}\right)\right)=A\cdot(1-0)=A\\[0.3cm]
\boxed{v_\infty=A}"

Now we can conclude about the physical meaning of the constant "A" : "A" is the boundary speed for a given type of motion.


ANSWER



"\left\{\begin{array}{l}
[A]=\left[\displaystyle\frac{m}{sec}\right]\\[0.3cm]
A-\text{the boundary speed}\\[0.3cm]
v(0)=0\\[0.3cm]
v_\infty=A
\end{array}\right."

QUESTION 2


To plot the graph, I chose the following constants:



"A=10\\[0.3cm]
t_{maxspeed}=5"




QUESTION 3


As we know



"v(t)=\frac{dx(t)}{dt}\to x(t)=\int v(t)dx=\int \left(A\cdot\left(1-e^{-\displaystyle\frac{t}{t_{maxspeed}}}\right)\right)dt\\[0.3cm]
x(t)=A\cdot\left(t+t_{maxspeed}\cdot e^{-\displaystyle\frac{t}{t_{maxspeed}}}\right)+Const\\[0.3cm]
\boxed{x(t)=A\cdot t_{maxspeed}\cdot\left(\frac{t}{t_{maxspeed}}+e^{-\displaystyle\frac{t}{t_{maxspeed}}}\right)+Const}"

To find the starting position, we will assume that "Const=0" , then



"x(0)=A\cdot t_{maxspeed}\cdot\left(\frac{0}{t_{maxspeed}}+e^{-\displaystyle\frac{0}{t_{maxspeed}}}\right)\\[0.3cm]
\boxed{x(0)=A\cdot t_{maxspeed}}"

Using the same assumption "Const=0" , we can find the asymptote for "t\to+\infty" :



"x_\infty=\lim\limits_{t\to+\infty}\left(A\cdot t_{maxspeed}\cdot\left(\frac{t}{t_{maxspeed}}+e^{-\displaystyle\frac{t}{t_{maxspeed}}}\right)\right)=+\infty\\[0.3cm]
\boxed{x_\infty=+\infty}"

ANSWER


General equation of motion:


"x(t)=A\cdot t_{maxspeed}\cdot\left(\frac{t}{t_{maxspeed}}+e^{-\displaystyle\frac{t}{t_{maxspeed}}}\right)+Const"

Under the assumption that "Const=0" :



"\left\{\begin{array}{l}
x(0)=A\cdot t_{maxspeed}\\[0.3cm]
x_\infty=+\infty
\end{array}\right."

QUESTION 4


To plot the graph, I chose the following constants:



"A=10\\[0.3cm]
t_{maxspeed}=5"




QUESTION 5


As we know



"a(t)=\frac{dv(t)}{dt}=\frac{d}{dt}\left(A\cdot\left(1-e^{-\displaystyle\frac{t}{t_{maxspeed}}}\right)\right)\\[0.3cm]
\boxed{a(t)=\frac{A}{t_{maxspeed}}\cdot e^{-\displaystyle\frac{t}{t_{maxspeed}}}}"

Initial acceleration is



"a(0)=\frac{A}{t_{maxspeed}}\cdot e^{-\displaystyle\frac{0}{t_{maxspeed}}}\\[0.3cm]
\boxed{a(0)=\frac{A}{t_{maxspeed}}}"

To find the asymptote as "t\to+\infty" we calculate the limit :



"a_\infty=\lim\limits_{t\to+\infty}\left(\frac{A}{t_{maxspeed}}\cdot e^{-\displaystyle\frac{t}{t_{maxspeed}}}\right)=0\\[0.3cm]
\boxed{a_\infty=0}"

ANSWER



"\left\{\begin{array}{l}
a(t)=\displaystyle\frac{A}{t_{maxspeed}}\cdot e^{-\displaystyle\frac{t}{t_{maxspeed}}}\\[0.3cm]
a(0)=\displaystyle\frac{A}{t_{maxspeed}}\\[0.3cm]
a_\infty=0
\end{array}\right."

QUESTION 6


To plot the graph, I chose the following constants:



"A=10\\[0.3cm]
t_{maxspeed}=5"




QUESTION 7


Initial conditions "x(0)=400" and "v(0)=28" .

These initial conditions cannot be, since in point 2 we theoretically proved that the initial speed must be 0. Therefore, I can not answer this question. Moreover, you need to specify the value "t_{maxspeed}" , although from the same paragraph 2 we can conclude that "t_{maxspeed}=+\infty" , which is clearly not suitable for real tasks